### v1

# Non-autoregressive Generative Models for Reranking Recommendation

Abstract.
Contemporary recommendation systems are designed to address the daily requirements of users by delivering a curated list of items that align with the user’s specific demands or interests. In a multi-stage recommendation system, reranking plays a crucial role by modeling the intra-list correlations among items.
The key challenge of reranking lies in the exploration of optimal sequences within the combinatorial space of permutations. Recent research proposes a generator-evaluator learning paradigm, where the generator generates multiple feasible sequences and the evaluator picks out the best sequence based on the estimated listwise score. Generator is of vital importance, and generative models are well-suited for the generator function. Current generative models employ an autoregressive strategy for sequence generation. However, deploying autoregressive models in real-time industrial systems is challenging. Firstly, the generator can only generate the target items one by one and hence suffers from slow inference. Secondly, the discrepancy between training and inference brings an error accumulation. Lastly, the left-to-right generation neglects information from succeeding items, resulting in suboptimal performance.
Hence, we propose a Non-AutoRegressive generative model for reranking Recommendation (NAR4Rec) designed to enhance efficiency and effectiveness. To address challenges related to sparse training samples and dynamic candidates impacting model convergence, we introduce a matching model. Considering the diverse nature of user feedback, we propose a sequence-level unlikelihood training objective to distinguish feasible from unfeasible sequences. Additionally, to overcome the lack of dependency modeling in non-autoregressive models regarding target items, we introduce contrastive decoding to capture correlations among these items.
Extensive offline experiments on publicly available datasets validate the superior performance of our proposed approach compared to the existing state-of-the-art reranking methods. Furthermore, our method has been fully deployed in a popular video app Kuaishou with over 300 million daily active users, significantly enhancing online recommendation quality, and demonstrating the effectiveness and efficiency of our approach.
recommender systems, generative model, non-autoregressive models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methods  Generative models
1. Introduction
Recommendation systems provide users with a personalized list of items that are tailored to their needs by aligning the recommendation with the user’s interest. With the emergence of deep learning, various approaches have been proposed to discern and capture user interests through the analysis of their actions. These models primarily focus on intricate feature interaction(Cheng et al., 2016; Guo et al., 2017; Lian et al., 2018), user preference modeling(Zhou et al., 2018, 2019b) and so on. However, most existing methods focus on modeling individual items separately using an itemwise learning paradigm. However, those methods fail to capture the mutual influence between items, resulting in suboptimal performance. Acknowledging that whether the user interactions with one item may correlate with others in the recommendation list(Pei et al., 2019), the reranking stage is introduced to consider the context information and then generate an optimal sequence of items of recommended items.
The key challenge of reranking lies in the exploration of optimal sequences within the combinatorial space of permutations. Existing methods for reranking can be categorized into one-stage and two-stage methods. One-stage methods(Ai et al., 2018; Pei et al., 2019) take candidates as input, estimating refined scores for each item under the permutation, and then rerank the candidates based on these refined scores in a greedy manner. However, one-stage methods encounter an inherent contradiction(Feng et al., 2021; Xi et al., 2021): the reranking operation inherently alters the permutation, introducing distinct mutual influences between items compared to the initial permutation. Consequently, the refined score conditioned on the initial permutation is considered implausible, indicating a coherence limitation in these one-stage methods.
To overcome this challenge, two-stage methods employ a generator-evaluator framework where the generator produces multiple feasible sequences, and the evaluator selects the optimal sequence based on the estimated listwise score. In the generator-evaluator framework, the generator is of vital importance. Generative models(Bello et al., 2018; Feng et al., 2021; Gong et al., 2022; Zhuang et al., 2018) are preferable over heuristic methods(Feng et al., 2021; Lin et al., 2023; Xi et al., 2021; Shi et al., 2023) for the generator function considering the expansive solution space of item permutations. Generative models utilized in this framework commonly employ an autoregressive strategy for sequence generation. However, deploying the autoregressive models in real-time industrial recommendation systems poses challenges.
Firstly, autoregressive models suffer from inference efficiency. Autoregressive models adopt a sequential approach to generate target sequences item by item, resulting in slow inference. The time complexity scales linearly with the sequence length during generation, leading to high latency in real-time applications.
Secondly, a critical issue arises from the training-inference discrepancy in autoregressive models. While these models are trained to predict the next item based on the ground truth up to that timestep, during inference, they receive their own previously generated outputs as input. This misalignment introduces an accumulated error, where inaccuracies generated in earlier timesteps propagate and accumulate over time. Consequently, this accumulation leads to sequences that deviate from the true distribution of the target sequence.
Additionally, autoregressive models exhibit limited information consideration. The sequential decoding process focuses solely on preceding items, neglecting information from succeeding items. This limitation results in suboptimal performance as the model fails to fully leverage the available context.
Figure 1. Comparison of autoregressive model (left) and non-autoregressive model (right). Auto-regressive models generate items sequentially while non-autoregressive generate all items simultaneously.
To address those challenges, we propose Non-AutoRegressive generative model for reranking Recommendation, referred to as NAR4Rec. As illustrated in fig. 1, traditional autoregressive models adopt a step-by-step approach to generate target sequences, where the model takes its own previously generated outputs as input. Conversely, our proposed NAR4Rec model generates all items within the target sequence simultaneously in a single model forward pass.
However, we find it nontrivial to deploy non-autoregressive in recommendation systems.
  • Firstly, the sparse nature and dynamic candidates in recommendation systems pose difficulties for learning convergence, which we address by sharing position embedding and introducing a matching model.
  • Secondly, the diverse nature of user feedback, including both positive and negative interactions, renders maximum likelihood training less suitable. We propose unlikelihood training to distinguish between desirable and undesirable sequences.
  • Lastly, canonical non-autoregressive models assume independent selection of items at each position in a sequence, which is inadequate when modeling intra-list correlation. We propose contrastive decoding to capture dependencies across items and address this limitation.
To summarize, our contributions are listed as follows:
  • To the best of our knowledge, we make the first attempt to adopt non-autoregressive models for reranking, which speeds up the inference speed and meets the requirements of real-time recommendation systems.
  • We propose a matching model for better convergence, a sequence-level unlikelihood training method for guiding the generated sequence towards better overall utility, and a contrastive search to constrain current decoding strategies with intra-list correlation.
  • Extensive offline experiments conducted on public dataset demonstrate that NAR4Rec outperforms state-of-the-art methods. Online experiments also shows the effectiveness of NAR4Rec. Furthermore, NAR4Rec have fully deployed in a real-world video app with over 300 million daily active users, significantly improve the recommendation quality.
2. Related work
2.1. Reranking in Recommendation Systems
In contrast to earlier phases like matching and ranking(Burges et al., 2005; Liu et al., 2009), which typically learn a user-specific itemwise scoring function, the core of reranking in recommendation systems lies in modeling correlations within the exposed list. Reranking(Pang et al., 2020; Pei et al., 2019; Xi et al., 2021), building upon candidate items from the ranking stage, selects a subset and determines their order to ensure exposing the most suitable items to the users. Existing research on reranking can be systematically classified into two principal categories: one-stage(Ai et al., 2018; Pei et al., 2019; Pang et al., 2020), and two-stage methods(Feng et al., 2021; Shi et al., 2023; Lin et al., 2023; Xi et al., 2021).
One-stage methods treat reranking as a retrieval task, recommending the top k items based on scores from a ranking model. These methods refine the initial list distribution using list-wise information, optimizing overall recommendation quality. Subsequently, the candidates are reranked by the refined itemwise score in a greedy manner. The distinction lies in the network architectures for capturing list-wise information, such as GRU in DLCM(Ai et al., 2018), and transformer in PRM(Pei et al., 2019). However, user feedback for the exposed list is influenced not just by item interest but also by arrangements and surrounding context(Joachims et al., 2017; Lorigo et al., 2008, 2006; Yang, 2017). The reranking operation modifies permutations, thereby introducing influences distinct from the initial permutation. Moreover, ne-stage methods, which exclusively model the initial permutation, fall short in capturing alternative permutations. Consequently, those methods struggle to maximize overall user feedback(Xi et al., 2021).
Two-stage methods (Feng et al., 2021; Shi et al., 2023; Lin et al., 2023; Xi et al., 2021) embrace a generator-evaluator framework. In this framework, the generator initiates the process by generating multiple feasible sequences, and subsequently, the evaluator selects the optimal sequence based on the estimated listwise score. This framework allows for a comprehensive exploration of various feasible sequences, and an informed selection of the most optimal one based on listwise considerations. The role of the generator is particularly crucial for generating sequences. Common approaches for generator can be categorized into heuristic methods(Feng et al., 2021; Lin et al., 2023; Xi et al., 2021; Shi et al., 2023), such as beam search or item swapping, and generative models (Bello et al., 2018; Feng et al., 2021; Gong et al., 2022; Zhuang et al., 2018). Generative models are more suitable than heuristic methods for reranking given the vast permutation space. These generative models typically adopt a step-greedy strategy which autoregressively decides the display results of each position. However, the high computational complexity of online inferences limits their application in real-time recommendation systems.
To address the challenges associated with autoregressive generation models, our work explores the potential of non-autoregressive generative models as the generator in the generator-evaluator framework. Non-autoregressive generative models generate the target sequence once to alleviate computational complexity.
2.2. Non-autoregressive Sequence Generation
Non-autoregressive sequence generation(Gu et al., 2018) is initially introduced in machine translation to speed up the decoding process. Then non-autoregressive has gained increased attention in nature language processing, e.g. text summarization(Gu et al., 2019; Awasthi et al., 2019), text error correction(Leng et al., 2021b, a). In particular, research efforts have been directed toward addressing challenges arising from the lack of target information in non-autoregressive models. Strategies include enhancing the training corpus to mitigate target-side dependencies(Gu et al., 2018; Zhou et al., 2019a) and refining training approaches(Ghazvininejad et al., 2019; Stern et al., 2019) to alleviate learning difficulties.
Though non-autoregressive generation has been explored in Natural Language Processing, the conventional techniques are not directly applicable to recommendation systems. We tackle the challenges encountered by recommendation to improve the convergence and performance of non-autoregressive models and make the first attempt to integrate non-autoregressive models into reranking in real-time recommender systems.
3. Preliminary
3.1. Reranking problem Formulation
For each user $u$ within the set $U$, a request encompasses a set of user profile features(such as user ID, gender, age), the recent interaction history and $n$ candidates items denoted as $X=\left\{x_{1},x_{2},\cdots,x_{n}\right\}$, where $n$ is the number of candidates. Given candidates $X$, the goal of reranking is to propose an item sequence $Y=\left\{y_{1},y_{2},\cdots,y_{m}\right\}$ with the most favorable feedback for user $u$, where $m$ is the sequences length. We denote the reranking models as $\mathcal{F}(u,\theta,X)$ where the corresponding parameter is $\theta$ and $Y$ is the outcome of the reranking model. In real-time recommendation systems, reranking acts as the last stage to deliver the ultimate list of recommended items. Typically, $n$ significantly exceeds $m$, with $m$ being less than 10 and $n$ ranging from several tens to hundreds.
In a multi-interaction scenario, users may exhibit distinct types of interaction (e.g., clicks, likes, comments) for each item exposure. Formally, we define the set of user interactions as $B$, and $y_{u,i,b}$ represents user $u$’s response to item $y_{i}$ concerning interaction $b\in B$. Given ${Y}$, each item $y_{i}$ elicits a multi-interaction response $\mathbf{s}_{u,y_{i}}=[{s}_{u,y_{i},b_{1}},\ldots,{s}_{u,y_{i},b_{|{B}|}}]$ and the overall user response is:
(1) $\mathbf{S}_{u,Y}=\left[\begin{matrix}s_{u,y_{1},b_{1}}&\dots&s_{u,y_{m},b_{1}}\\ \vdots&\ddots&\vdots\\ s_{u,y_{1},b_{|B|}}&\dots&s_{u,y_{m},b_{|B|}}\end{matrix}\right].$
The overall utility is quantified as the summation of individual item utilities, denoted as $\mathcal{R}(u,Y)=\sum_{i=1}^{m}\mathcal{R}(u,y_{i})$. The utility associated with each item may correspond to a specific interaction type $b$, such as clicks, watch time, or likes. In such cases, the item utility is expressed as $\mathcal{R}(u,y_{i})=s_{u,y_{i},b}$. Alternatively, the item utility can be represented as the weighted sum of diverse interactions $\mathcal{R}(u,y_{i})=\sum_{b}w_{b}s_{u,y_{i},b}$, where $w_{b}$ denotes the weight for each interaction.
The reranking objective is to maximum the overall utility $\mathcal{R}(u,Y)$ for a given user $u$:
(2) $max_{\theta}\mathcal{R}(u,Y).$
The reranking phase introduces a permutation space with exponential size, precisely expressed as $\mathcal{O}(A_{n}^{m})$. This denotes the solution space for selecting and ordering $m$ items from a pool of $n$ candidates. Considering that items in a list have intricate connections to each other, users display unique feedback for each permutation. However, in practical scenarios, users are exposed to only one permutation. Consequently, the primary challenge in reranking lies in efficiently and effectively determining the permutation-optimal list given the vast solution space yet extremely sparse real user feedback as training samples.
3.2. Autoregressive sequence generation
Given a set of candidate items denoted as $X$, autoregressive models decompose the distribution over potential generated sequences $Y$ into a series of conditional probabilities:
(3) $p_{\mathcal{\text{AR}}}(Y|X;\theta)=\sum_{t=1}^{m+1}p(y_{t}|y_{0:t-1},x_{1:n};\theta).$
where the special tokens $y_{0}$ (e.g., ¡bos¿) and $y_{m+1}$ (e.g., ¡eos¿) denote the beginning and end of target sequences. Importantly, the length of the generated sequence is predetermined and fixed, unlike variable lengths in natural language generation.
Factorizing the sequence generation output distribution autoregressively leads to a maximum likelihood training with a cross-entropy loss at each timestep:
(4) $\mathcal{L}_{\text{AR}}=-\log p_{\text{AR}}(Y|X;\theta)=-\sum_{t=1}^{m+1}\log p(y_{t}|y_{0:t-1},x_{1:n};\theta).$
The training objective aims to optimize individual conditional probabilities. During training, where the target sequence is known, the computation of these probabilities is conditioned on the target items at earlier timesteps rather than the model-generated output items, allowing for efficient parallelization. During inference, the autoregressive models generate the target sequence item-by-item sequentially. Autoregressive models align well with the inherent browsing nature of recommendation systems, efficiently capturing the distribution of the target sequence. This makes them well-suited for the reranking task, particularly considering the vast space of possible permutations.
While autoregressive models have demonstrated effectiveness, their deployment in industrial recommendation systems is challenging. Firstly, their sequential decoding nature leads to slow inference, introducing latency that hinders real-time application. Secondly, these models, trained to predict based on ground truth, face a discrepancy during inference, receiving their own generated outputs as input. This misalignment may lead to compounded errors, as inaccuracies generated in earlier timesteps accumulate over time, resulting in inconsistent or divergent sequences that deviate from the true distribution of the target sequence. Thirdly, autoregressive models rely on a left-to-right causal attention mechanism, limiting the expressive power of hidden representations, as each item encodes information solely from preceding items(Sun et al., 2019). This constraint impedes optimal representation learning, resulting in suboptimal performance.
3.3. Non-autoregressive sequence generation
To address the challenges mentioned earlier, non-autoregressive sequence generation takes a new approach by removing autoregressive dependencies from existing models Each element’s distribution $p(y_{t})$ depends solely on the candidates $X$:
(5) $p_{\text{NAR}}(Y|X;\theta)=\prod_{t=1}^{m}p(y_{t}|x_{1:n};\theta).$
Then the loss function for the non-autoregressive model is:
(6) $\mathcal{L}_{\text{NAR}}=-\log p_{\text{NAR}}(Y|X;\theta)=-\prod_{t=1}^{m}\log p(y_{t}|x_{1:n};\theta).$
Despite the removal of the autoregressive structure, the models retain an explicit likelihood function. The training of models employs separate cross-entropy losses for each output distribution. Crucially, these distributions can be computed concurrently during inference, introducing a marked departure from the sequential nature of autoregressive decoding. This non-autoregressive approach mitigates inference latency, thereby enhancing the efficiency of recommendation systems in real-world applications.
4. Approach
We will first introduce our non-autoregressive model structure in detail, and then detail the techniques we proposed to solve the challenges to apply it in reranking, including unlikelihood training, contrastive decoding, and sequence evaluation.
4.1. Matching model
Non-autoregressive models face challenges in converging during training due to two main reasons. Firstly, the sparse nature of training sequences makes learning difficult. Unlike text sequences that often share linguistic structures, recommended lists in training samples rarely have the same exposures, making effective learning from limited data challenging. Secondly, in the reranking stage, the same index for candidates can represent different items, resulting in a variable vocabulary as the candidates to be ranked vary across samples. Conventional models may struggle to handle such variations efficiently. To tackle these challenges, we introduce two key components to our models: a candidates encoder for effectively encoding representations of candidates and a position encoder to capture position-specific information within the target sequence.
To start, we randomly initialize an embedding for each position in the target sequences. Notably, we share these position embeddings across training data to enhance learning on sparse data. We then incorporate bidirectional self-attention and cross-attention modules to learn representations for each position, leveraging information from the candidates.
Additionally, to address the challenge posed by the dynamic vocabulary arising from variations in candidates across samples, we employ a matching mechanism. Specifically, we match the representation of each candidate with each position in the target sequence, resulting in the computation of probabilities for each candidate at every position.
Taking a user and the corresponding list of candidate items as input, our designed model aims to output a probability matrix $\hat{P}\in\mathbb{R}^{n\times m}$, where $\hat{P}_{i,j}$ indicates the matching score of placing the $i$-th item in the $j$-th position of the final list. Specifically, the model consists of a profile embedding layer, a candidates encoder, and a position encoder. The input profiles are first sent to the profile embedding layer and embedded into the initial candidates embeddings. Then, the candidates encoder further processes these initial embeddings via a self-attention mechanism to capture the mutual influence between items. Alongside this, the position encoder also processes a set of learnable position vectors in the same way. Finally, the output embeddings of the candidates encoder and the position encoder are used to compute the score matrix by matrix multiplication.
In the following, we describe the structure design of each model component in detail.
Embedding layer. Following previous works, we parameterize the available profiles into vector representations for users and items. Given a user $u$, associated with sparse features $x_{u}^{s}$ and dense features $x_{u}^{d}$, we embed each sparse feature value into $d$-dimensional space, while handling dense features standardization or batch normalization to ensure normal distribution. Subsequently, each user $u$ can be represented as $x_{u}\in\mathbb{R}^{|x_{u}^{s}|\times d+|x_{u}^{d}|}$, where $|x_{u}^{s}|$ and $|x_{u}^{d}|$ denote the size of sparse and dense feature space of user $u$ respectively. Similarly, we represent each item $i$ as $x_{i}\in\mathbb{R}^{|X_{i}^{s}|\times d+|x_{i}^{d}|}$. Then the representation for item candidates is denoted as $X=\{x_{1},x_{2},\cdots,x_{n}\}$
Also, we randomly initialize an embedding for each position $j$ as $p_{j}\in\mathbb{R}^{d}$. Note that these embeddings are learnable during the training process, and are fixed when using the model to evaluate the quality of the candidate list. The representation for positions is denoted as $P\in\mathbb{R}^{m\times d}$.
Candidates Encoder. The candidates encoder adopts the standard Transformer architecture by stacking $N$ Transformer layers. In each layer, the architecture mainly consists of a self-attention block and a feed forward network. Denoted the input feature as $\mathbf{X}$, the self-attention block first transform it into query ($\mathbf{Q}$), key ($\mathbf{K}$) and value ($\mathbf{V}$) matrices via linear projection, as follows:
(7) $\mathbf{Q}=\mathbf{X}\mathbf{W}^{\mathbf{Q}},\mathbf{K}=\mathbf{X}\mathbf{W}^{\mathbf{K}},\mathbf{V}=\mathbf{X}\mathbf{W}^{\mathbf{V}},$
where $\mathbf{X^{Q}}$, $\mathbf{X^{K}}$ and $\mathbf{X^{V}}$ denote the weight matrices. Then, the self-attention operation is applied as:
(8) $Attention(\mathbf{Q},\mathbf{K},\mathbf{V})={\rm Softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}})\mathbf{V},$
where $d_{k}$ denotes the dimension of the $\mathbf{K}$ matrix. Finally, it applies a linear projection to obtain the output of the self-attention block.
The feed forward network is typically placed after the self-attention block. It consists of two linear transformation layers and a non-linear activation function. It receives the output features $\mathbf{X}$ from the self-attention block, and processes them as:
(9) $FFN(\mathbf{X})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{X}),$
where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ denote the weight matrices of the two linear layers.
Position Encoder. The position encoder adopts a similar Transformer architecture as the candidates encoder. The only difference between them is that the position encoder inserts a cross attention block between the self-attention block and the feed forward network in each Transformer layer. As can be seen in Fig. 2, in each layer, the cross attention block receives the output features from the self-attention blocks in the same layer of both encoders and processes them via cross attention operation. Specifically, denoted $\mathbf{X}^{\mathbf{CE}}$ as the feature from the candidates encoder, and $\mathbf{X}^{\mathbf{PE}}$ as the feature from the position encoder, similar to the self-attention block, it first applies the linear projection to them:
(10) $\mathbf{Q}=\mathbf{X}^{\mathbf{CE}}\mathbf{W}^{\mathbf{Q}},\mathbf{K}=\mathbf{X}^{\mathbf{PE}}\mathbf{W}^{\mathbf{K}},\mathbf{V}=\mathbf{X}^{\mathbf{PE}}\mathbf{W}^{\mathbf{V}}.$
Then, it applies the formula in Eq. 8 to $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ to obtain the output feature. The cross attention uses the features from the candidates encoder as query and features from the position encoder as key and value, which can establish a connection between the embeddings of candidate items and the position.
Probability Matrix. Given the output features of the candidates encoder (denoted as $\{\mathbf{x_{1}},\mathbf{x_{2}},...,\mathbf{x_{n}}\}$) and position encoder (denoted as $\{\mathbf{p_{1}},\mathbf{p_{2}},...,\mathbf{p_{m}}\}$), we apply a matrix multiplication to them to compute the matching scores. A column-wise softmax function is then adopted to further normalize the scores. Formally, the matching score of placing the $i$-th candidate item to the $j$-th position is computed as:
(11) $\hat{p_{ij}}=\frac{exp(\mathbf{x_{i}}^{\intercal}\mathbf{p_{j}})}{\sum_{i=1}^{n}exp(\mathbf{x_{i}}^{\intercal}\mathbf{p_{j}})}$
where $p_{ij}$ is 1 if $x_{j}$ is in position $p_{j}$ otherwise 0.
Training objectives Our model is optimized via cross-entropy loss function, which is defined as follows:
(12) $\mathcal{L}(Y,X)=-\sum_{i=1}^{n}\sum_{j=1}^{m}p_{ij}\log(\hat{p_{ij}})$
Figure 2. Architectural Overview of the Generator and Evaluator Models. The evaluator evaluates multiple sequences generated by the generator and estimates listwise scores to select the optimal sequence.
4.2. Unlikelihood Training
The divergence between text and item sequences hinders the direct application of generative models from natural language processing for item sequence recommendation. This distinction originates from the distinctive features of user interactions in recommendation scenarios. The varied feedback within recommendation sequences differs from the structured nature of natural language text. While text sequences typically adhere to standard language structures, conveying information or forming a coherent narrative, user feedback in recommendation sequences is diverse, encompassing both positive and negative interactions. Users manifest preferences through actions like clicks or likes, contributing to a diverse and nuanced feedback landscape.
Hence, the difference in objectives between maximum likelihood training(as in eq. 5) and reranking (as in eq. 2) poses a significant challenge. Although maximum likelihood training effectively captures patterns in text sequences, its applicability diminishes in recommendation scenarios where user preferences are dynamic and subjective. The essence of high-quality recommendations lies not just in sequence patterns from training data but, more crucially, in the user utility of the recommended list. User interactions with recommended items are subjective and context-dependent, introducing complexity in aligning the training goal with the desired model behavior. To navigate this challenge, we propose unlikelihood training to guide the model to assign lower probabilities to undesirable generations. This adjustment aligns the training process with the intricate feedback patterns inherent in recommendation scenarios.
Unlikelihood training decreases the model’s probability of generating a negative sequence. Given candidates $X$ and a negative sequence $Y$, the unlikelihood loss is:
(13) $\mathcal{L}_{\text{ul}}(Y_{neg},X)=-\sum_{i=1}^{n}\sum_{j=1}^{m}p_{ij}\log(1-\hat{p}_{ij})$
The loss decreases as $\hat{p}_{ij}$ decreases.
Unlike regular text generation, where messages are clear and content-focused, the manageability of attributes like topic, style, and sentiment in the output text is evident and systematically controlled(Li et al., 2020; Welleck et al., 2019). However, recommendation sequences involve user feedback with implicit signals. For instance, a lack of interaction with a recommended item may suggest disinterest. This highlights the model’s need to understand both explicit and implicit cues in user feedback. Effective control over a generation in recommendation sequences becomes crucial to tailor the output based on users’ preferences and behaviors, thus enhancing the personalized nature of recommendations. Specifically, we regard a sequence as a positive or negative sequence based on the overall utility defined in section 3.1, and the corresponding loss is as follows:
(14) $\mathcal{L}_{\text{ul}}(Y,X)=\left\{\begin{array}[]{ll}-\sum_{i=1}^{n}\sum_{j=1}^{m}p_{ij}\log(1-\hat{p}_{ij})&\text{ if }\mathcal{R}(u,Y)<\alpha\\ -\sum_{i=1}^{n}\sum_{j=1}^{m}P_{ij}\log(\hat{P_{ij}})&\text{ if }\mathcal{R}(u,Y)\geq\alpha\end{array}\right.$
where $\alpha$ is the threshold for positive and negative sequences.
In summary, beyond the primary goal of learning positive sequence patterns through sequence likelihood, unlikelihood training introduces an additional objective to reduce the likelihood of generating sequences with low utilities, effectively training recommendation models to discern feedback within recommendation sequences.
4.3. Contrastive Decoding
Compared with autoregressive generation, the non-autoregressive approach significantly enhances computation efficiency and makes it feasible to deploy in real-time recommendation systems. However, non-autoregressive generation uses the conditional independence assumption: each target item’s distribution $p(y_{t})$ depends only on the candidates $X$. This deviation from autoregressive models challenges in capturing the inherently multimodal nature of the distribution of valid target sequences. Take machine translation for example, when translating the phrase ”thank you” into German could result in multiple valid translations such as ”Vielen Dank” and ”Danke”. However, non-autoregressive models may generate unplausible translations like ”Danke Dank”. The conditional independence assumption in eq. 5 restricts the model’s ability to effectively capture the multimodal distribution in the set of possible target sequences.
Essentially, the assumption of conditional independence limits the model’s ability to navigate a vast solution space and identify the most suitable permutation from numerous valid options for a given set of candidates. This limitation is especially notable in scenarios where the number of reasonable target sequences far exceeds those encountered in natural language processing. Consequently, non-autoregressive frameworks face the challenge of addressing the impact of conditional independence to improve their capacity for generating diverse and contextually appropriate target sequences. To tackle this, we propose contrastive decoding to model the co-occurrence relationship between items and thereby improve the target dependency.
Contrastive decoding incorporates a diversity prior that regulates the sequence decoding procedure with the predicted distribution of the model. This is based on the intuition that a good recommended list needs to be composed of a wide variety of items. In fact, contrastive decoding leverages a similarity score function as a regularizer when decoding, which captures the dependence between different positions in the target sequence.
Formally, given the previous context $y_{<t}$, at time step $t$, the selection of the output $y_{t}$ follows
(15) $y_{t}=argmax_{x\in X}(1-\alpha)\times p(x|p_{i},X)-\alpha\times max(c(\mathbf{x}_{i},\mathbf{x}_{j}))$
where $11\leq j\leq t-$ 1. In eq. 15, the first term, model confidence is the probability of candidates $x$ predicted by the model. The second term, similarity penalty, measures how discriminative of candidate $x$ with respect to the previously selected items, where $c(\mathbf{x_{i}},\mathbf{x_{j}})$
(16) $c(\mathbf{x_{i}},\mathbf{x_{j}})=\frac{\mathbf{x_{i}}^{\top}\mathbf{x_{j}}}{\|\mathbf{x_{i}}\|\cdot\|\mathbf{x_{j}}\|}.$
Specifically, the similarity penalty is defined as the maximum similarity between the representation of $x$ and all items in $y_{<t}$. Intuitively, a larger similarity penalty it is more similar to the context.
Also, to encourage the language model to learn discriminative and isotropic item representations, we incorporate a contrastive objective into the training of the language model. Specifically, given a sequence $X$, the $\mathcal{L}_{\textup{CL}}$ is defined as
(17) $\mathcal{L}_{\text{position}}=\frac{1}{n\times(n-1)}\sum_{i=1}^{n}\sum_{j=1,j\neq i}^{n}\max\{0,\rho-c(\mathbf{x}_{i},\mathbf{x}_{i})+c(\mathbf{x}_{i},\mathbf{x}_{j})\},$
where $\rho\in[-1,1]$ is a pre-defined margin and $x_{i}$ is the representation of item $x_{i}$ produced by the model.
(18) $\mathcal{L}_{\text{item}}=\frac{1}{m\times(m-1)}\sum_{i=1}^{m}\sum_{j=1,j\neq i}^{|\boldsymbol{x}|}\max\{0,\rho-c(\mathbf{p}_{i},\mathbf{p}_{i})+c(\mathbf{p}_{i},\mathbf{p}_{j})\},$
Intuitively, by training with $\mathcal{L}_{\textup{CL}}$, the model learns to pull away the distances between representations of distinct tokens. [^1] By definition, the cosine similarity $s(h_{x_{i}},h_{x_{i}})$ of the identical token $x_{i}$ is $1.0$. Therefore, a discriminative and isotropic model representation space can be obtained. The overall training objective $\mathcal{L}$ is then defined as
(19) $\mathcal{L}(Y,X)=\mathcal{L}_{\text{ul}}+\mathcal{L}_{\text{position}}+\mathcal{L}_{\text{item}},$
where the maximum likelihood estimation (MLE) objective $\mathcal{L}$ is described in eq. 6. Note that, when the margin $\rho$ in $\mathcal{L}_{\textup{CL}}$ equals to 00, the $\mathcal{L}_{\textup{SimCTG}}$ degenerates to the vanilla MLE objective $\mathcal{L}_{\textup{MLE}}$.
4.4. Sequence Evaluator
The sequence evaluator model is designed to estimate the overall utility of a given sequence, as illustrated in fig. 2. The generated sequence from the generator undergoes encoding through self-attention and a feed-forward layer to obtain contextual information. The positional representation passes the linear projection layer to predict the score for a specific target. The overall utility is calculated as the weighted sum of scores at each position.
5. Experiments
We conduct extensive experiments in both offline and online environments to demonstrate the effectiveness of NAR4Rec. Three questions are investigated in the experiments:
  • How does NAR4Rec perform in comparison with state-of-the-art performance in terms of recommendation quality?
  • How do different settings of NAR4Rec affect the performance?
  • How does NAR4Rec perform in real-life recommendation systems?
5.1. Experiment details
Dataset: To ensure dataset consistency and align with the problem setup, we emphasize the inclusion of real item sequences displayed in complete sessions, rather than manually constructed lists. Consequently, we have collected two datasets specifically for experiments.
  • Avito Dataset: The Avito dataset is a publicly available collection of user search logs. Each sample represents a search page containing multiple ads, forming a really impressive list with feedback. The dataset comprises over 53 million lists involving 1.3 million users and 36 million ads. The training phase utilizes impressions from the first 21 days, while the testing phase utilizes impressions from the last 7 days. Each list in the dataset has a length of 5.
  • VideoRerank Dataset: The VideoRerank dataset is derived from Kuaishou, a widely used short-video application with a user base of over 300 million daily active users. Each sample in the dataset represents an actual sequence of items displayed to a specific user during a complete session, ensuring dataset consistency and problem alignment. The dataset consists of a total of 82,230,788 users, 26,835,337 items, and 1,811,625,438 requests. Each request includes 6 items and 60 candidates, about the ranking process.
Baselines We conduct a comparative analysis of our NAR4Rec approach with five widely adopted reranking methods, namely DNN, DCN, PRM, PIER, and Edge-Rerank, which serve as our baselines. Here is a brief overview of these baseline methods:
  • DNN(Covington et al., 2016): DNN is a fundamental method for click-through rate (CTR) prediction that utilizes a multi-layer perceptron (MLP) to capture high-order feature interactions.
  • DCN(Wang et al., 2017): DCN incorporates feature crossing at each layer, eliminating the need for manual feature engineering while keeping the added complexity to the DNN model minimal.
  • PRM(Pei et al., 2019): PRM enhances an initial list by leveraging the self-attention mechanism to account for the mutual influence among items.
  • Edge-Rerank(Gong et al., 2022): Edge-Rerank combines an on-device ranking model with an adaptive beam search approach to generate context-aware re-ranking outcomes.
  • PIER(Shi et al., 2023): PIER treats the target ad and all slots as a query and employs external items as key-value pairs to model externalities in different exposure situations.
Metrics In our experiments, we employ commonly used metrics to evaluate the performance of different methods on the Avito dataset and the VideoRerank dataset. Specifically, we utilize AUC and LogLoss as evaluation metrics for the Avito dataset, while for the VideoRerank dataset, which contains both candidates and exposed item lists, we employ Recall@6, Recall@10, and LogLoss as evaluation metrics.
5.2. Offline experiments
In this section, we conducted offline experiments to evaluate the effectiveness of our method on two datasets.
Hyperparameters In our offline experiments with the Avito dataset, we use the following hyperparameters: a learning rate of 0.001, a batch size of 1024, and the Adam optimizer. For the VideoRerank dataset, we maintain a learning rate of 0.001, a batch size of 256, and utilize the Adam optimizer as well.
5.2.1. Performance comparison(RQ1)
The experimental results on the Avito dataset are presented in table 1, following the approach proposed by (Shi et al., 2023). AUC and LogLoss are the evaluation metrics used in this study. Our method achieves the best performance among the baseline models in both AUC and LogLoss. PRM outperforms DNN and DCN by effectively capturing the mutual influence between items. Additionally, Edge-Rerank surpasses DNN and DCN by integrating an on-device ranking model with an adaptive beam search technique. PIER demonstrates superiority over other baselines by considering the interaction per category. Notably, our proposed method exhibits the highest improvement with a significant increase of 0.0125 in the AUC metric compared to other baseline models.
Table 1. Comparison between Nar4Rec and other baselines on Avito.
AUC LogLoss
DNN 0.6614 0.0598
DCN 0.6623 0.0598
PRM 0.6881 0.0594
Edge-rerank 0.6953 0.0574
PIER 0.7109 0.0409
NAR4Rec 0.7234 0.0384
5.2.2. Analysis of NAR4Rec(RQ2)
As shown in fig. 3 and fig. 4, by conducting a series of experiments on NAR4Rec, we demonstrate that our experimental results exhibit insensitivity to variations in the learning rate and batch size.
Figure 3. The performances of NAR4Rec on Avito Dataset with different learning rate.
Figure 4. The performances of NAR4Rec on Avito Dataset with different batch size.
Table 2 shows the results of our offline experiments on VideoRerank. The evaluation metrics used in this experiment include Recall@6, Recall@10, and Loss. Our method achieves superior results on all metrics compared to other baseline models as well.
Table 2. Comparison between Nar4Rec and other baselines on VideoRerank
Recall@6 Recall@10 LogLoss
DNN 66.47% 86.65% 0.6764
DCN 68.22% 87.95% 0.6809
PRM 73.17% 92.25% 0.5328
PIER 73.50% 92.44% 0.5361
Edge-rerank 73.63 % 92.90% 0.5252
NAR4Rec 74.86% 93.16% 0.5199
5.3. Online experiments(RQ3)
We conduct online A/B experiments on the KuaiShou App.
5.3.1. Experiments setup
In online A/B experiments, we evenly divided the traffic of the entire app into ten buckets. The online baseline is Edge-rerank(Gong et al., 2022), with 20% of the traffic assigned to NAR4Rec, while the remaining traffic continues with Edge-rerank. We set two conditions for NAR4Rec: first, we expect the users to watch more videos and stay more time on each video; second, we expect positive feedbacks from users.
5.3.2. Experimental Results
The experiments have been launched on the system for ten days, and the result is listed in table 3. NAR4Rec outperforms Edge-rerank(Gong et al., 2022) by a large margin.
5.3.3. Ablation study
Given the absence of an evaluation metric for the generated sequence, we conducted online experiments to demonstrate the effectiveness of NAR4Rec. The removal of the contrastive search resulted in a decrease in diversity by +0.847%. Furthermore, training on all exposed item lists led to a significant drop in watch time by -1.0% (considering a drop of around 0.2% as significant in our system).
Views Likes Follow Long Views Complete Views
+1.161% +1.71% +1.15% +1.82% +2.45%
Table 3. Online experiments results. All values are the relative improvements of NAR4Rec. For online A/B test in Kuaishou, the improvements of over 0.5% in position interactions(like, follow) and 0.2% in views are very significant.
6. Conclusion
In this paper, we have presented an overview of the current formulation and challenges associated with reranking in recommendation systems. Though non-autoregressive generation has been explored in natural language processing, the conventional techniques are not directly applicable to recommendation systems. We tackle the challenges encountered by recommendation to improve the convergence and performance of non-autoregressive models and make the first attempt to integrate non-autoregressive models into reranking in real-time recommender systems. Extensive online and offline experiments have demonstrated the efficacy of NAR4Rec as a versatile framework for generating sequences with enhanced utility. Moving forward, our future work will focus on refining the modeling of sequence utility to further enhance the capabilities of NAR4Rec.